Verbal Test 90 70 78 72 89 75 78 56 80 2. 60 92 82 98 Display each of these variables in a frequency distri- bution with columns for percentages and cumulative percentages 2.8 GER Twenty-five residents of a community for sen- ior citizens were asked to keep track of the number of hours they watched TV during the past week. 20 2 10 15 21 18 12 32 7 13 17 14 15 4 10 27 2 5 11 a. Conastruet a frecquelay be data b. What are the midpoints of the class intervals? e. Add columns to the table to display the percentage distribution, cumulative frequencies, and cumula- tive percentages. d. Write a paragraph summarizing this distribution of Statisti scores. 11 IS
Answer #1

(a)

Class Interval Mid point Frequency Percentage frequency Cumulative frequency Cumulative percentage
0 - 4 2 3 (3/25)X100=12% 3 (3/25)X100=12%
5 - 9 7 8 (8/25)X100=32% 11 (11/25)X100=44%
10- 14 12 6 (6/25)X100=24% 17 (17/25)X100=68%
15 - 19 17 4 (4/25)X100=16% 21 (21/25)X100=84%
20- 24 22 2 (2/25)X100=8% 23 (23/25)X100=92%
25 - 29 27 1 (1/25)X100=4% 24 (24/25)X100=96%
30 - 34 32 1 (1/25)X100=4% 25 (25/25)X100=100%
Total = 25 100%

Columns 1 , 2 and 3 give the frequency distribution.

(b)

The mid points of the class intervals are given in column 2

(c)

Percentage distribution is given in column 4.

Cumulative distribution is given in column 5.

Cumulative percentages are given in column 6.

(d)

The distribution of scores is right skewed (Positive skewness). The distribution is not symmetric as it does not showany kind of symmetry.It has a long right tail. The mean is tothe right of the peak.